# LLMs Can Compensate for Deficiencies in Visual Representations

### v1

@@ Abstract
Many vision-language models (VLMs) that prove very effective at a range of multimodal task, build on CLIP-based vision encoders, which are known to have various limitations. We investigate the hypothesis that the strong language backbone in VLMs compensates for possibly weak visual features by contextualizing or enriching them. Using three CLIP-based VLMs, we perform controlled self-attention ablations on a carefully designed probing task. Our findings show that despite known limitations, CLIP visual representations offer ready-to-read semantic information to the language decoder. However, in scenarios of reduced contextualization in the visual representations, the language decoder can largely compensate for the deficiency and recover performance. This suggests a dynamic division of labor in VLMs and motivates future architectures that offload more visual processing to the language decoder.
@@ 1 Introduction
Vision-language models (VLMs) have made remarkable progress in recent years, with systems like MAGMA (Eichenberg et al., 2021), BLIP-2 (Li et al., 2023), LLaVA Liu et al. (2023, 2024a), and Prismatic (Karamcheti et al., 2024) demonstrating strong performance for their time on key multimodal benchmarks. A common design choice across these models is the use of a frozen pretrained vision encoder, often based on CLIP Radford et al. (2021), paired with a pre-trained language model, which is finetuned to map visual features into text. [^1] The vision encoder and language decoder are commonly connected with a linear projection or a multi-layer perceptron, which has largely been shown to serve the technical role of mapping between dimensionalities rather than semantic spaces (Schwettmann et al., 2023a; Verma et al., 2024).
@@ 2 Related Work
Zooming in on the vision encoder itself, Gandelsman et al. (2024, 2025) decomposed CLIP’s visual embeddings into sparse text concepts and found that different attention heads specialize in distinct semantic properties (e.g., color, shape, location, etc,). Despite CLIP’s strong zero-shot classification performance, various studies have highlighted its limitations, ranging from poor compositional reasoning and limited sensitivity to fine visual distinctions (Wang et al., 2025; Shao et al., 2023; Lewis et al., 2024; Zhang et al., 2024b), to biases related to object quantity and size Abbasi et al. (2025).
@@
Figure 1: Overview of Object Part Identification. Given an object part (e.g., the tail of a cat), localized to a region of image patches through a segmentation mask, a representation is obtain from the VLM for all relevant patches, a probability distribution is induced through LogitLens, and an identifiability rank is extracted from each distribution for the relevant label (tail). The highest rank across patches indicates the overall part identifiability.
Together, these studies tell a nuanced story of the mechanisms behind building visual representations, with their perceived and real limitations, and translating those into text. Our work investigates how visual representations evolve through the layers of the language decoder, and to what extent their deficiencies are resolved in the process.
@@ 3 Object Part Identification (OPI)
OPI assumes the availability of patch-level masks which localize object parts in an input image. A single patch can contain multiple parts and a single part can span multiple patches. Probing the VLM consists in: (1) inducing a probability distribution from a VLM for every image patch, (2) extracting the rank of all part labels relevant to a patch, and (3) aggregating the ranks across all patches in the part-relevant. This results in one value per part which indicates how well the model identifies this part in this image.
@@
We derive the necessary annotations from the Pascal Panoptic Parts (Pascal-PP) dataset (de Geus et al., 2021; Meletis et al., 2020) which provides high-quality pixel-level segmentation annotations for 194 part classes spanning 16 object classes, e.g., dog-head, horse-tail, person-leg, etc. We focus on 7 out of the 16 object classes available in Pascal-PP, specifically animal objects. This and other measures taken to reduce noise are outlined in Appendix A. After filtering, we sample up to 100 images per class, where available.
Finally, we convert pixel-level segmentation masks into patch-level masks, effectively reducing the resolution of the masks. The final dataset contains 567 unique images and 2840 unique part regions. (see Table 1 for more details.)
@@
Prior to observing the rank of a label, the probability distribution is processed to aggregate all label aliases e.g., “leg”, “legs”, “_leg”, “_legs”, under a single token, leg, summing up all their probabilities. [^3] The label aliases are manually defined with reference to the VLM vocabulary. The rank of label $l$ is obtained from the probability distribution $p_{i}$ as follows:
@@ 4.1 Model Architecture and Details
Table 1: Summary of selected object classes with available parts. $N$: total number of images per class and $P$: total number of unique part regions per class.
We carry out experiments with LLaVA-1.5 7B and 13B , BakLLaVA 7B and TinyLLaVA (Zhou et al., 2024), all three of which use the CLIP ViT-L/14 image encoder with 24 layers (Radford et al., 2021). As decoder, LLaVA-1.5 7B/13B uses Vicuna 7B/13B (Chiang et al., 2023), BakLLaVA uses Mistral 7B v0.1 (Jiang et al., 2023), and TinyLLaVA uses TinyLLaMA (1.1B) (Zhang et al., 2024a). All language decoders are prompted using their standard template (“USER: <image>”).
This selection of models with different language backbones allows us to robustly study the role of visual input contextualization in LLMs in general. Indeed, our results show consistent trends across all VLMs, with reduced magnitude for the considerably smaller TinyLLaVA model, whose results we show Table 3 in the Appendix.
@@ 4.2 Establishing a Baseline
Using the method introduced in §3, we obtain OPI scores from all models. The scores, averaged across all parts per object (bird, cat, etc.) and across all objects (all), can be seen in Figure 2 (left). Average object identifiability scores are also included for reference (hatched bars), as induced from part regions. We observe that the OPI rate is significantly above chance for all three models, at or above 80% on average, and matches the object identifiability. This shows that the VLMs effectively map between visual concepts and their corresponding vocabulary tokens even on the low level of object parts. While it is expected that object information is often mentioned in text, object parts are mentioned far more rarely.
In Figure 2 (right), we see how identifiability progresses across the layers of the models, as a function of percentage of layers computed. OPI in LLaVA 13B progresses in two steps, with one surge at about 20% and another at 40% after which it nearly plateaus. LLaVA 7B shows a steep ascent up to about 60% of its layers, and BakLLaVA’s progression is stable throughout. Interestingly, all models show a drop in identifiability in the last two-three layers (with a subsequent recovery in the last layer for LLaVA 7B only). This LLM behavior might be linked to confidence calibration mechanisms, which aim to prioritize a specific set of output tokens, subject to strong language priors in addition to visual context.
@@ 4.3 Ablating Contextualization
Having established that LLaVA is highly successful in identifying object parts, next we inspect the source of this performance by ablating the contextualization of images patches in the visual encoder and the language decoder through Attention Knockout (AK) (Geva et al., 2023). We block the flow of information between part regions and the rest of the image, at different layers of LLaVA. This allows us to assess how contextualization plays into OPI across different layers of the encoder and decoder.
@@
Figure 12 presents the layer-wise identifiability results for all experimental configurations, with a per-object breakdown for LLaVA 13B and averaged results for all three models. Below, each configuration is discussed in turn.
@@
To assess the role of contextualization in the OPI task, we establish a floor of performance by applying attention knockout to both the vision encoder and the language decoder (Full AK). This leads to a sizable drop in OPI scores compared to the baseline (No AK), which limited recovery across the layers of the decoder. Despite this drop, the identifiability rate remains nonzero, indicating that some object parts can be identified to a high degree even is isolation from the object they belong to. In Figure 4, we show one such case: the leg of a person, whose identifiability drops a bit as a results of the full attention knockout, but stands apart from other parts in how little its identifiability changes. Overall, however, OPI proves to be a highly context-dependent task, as intended.
Trends are similar for BakLLaVA 7B and LLaVA 7B (last column in Figure 4), with the gap between the upper (No AK) bound and lower (Full AK) bound being smaller in comparison to LLaVA 13B, but still considerable.
@@
Based on the previous two AK configurations, we would expect that blocking self-attention in the vision encoder (AK in CLIP Encoder) should have an almost as detrimental effect as blocking all self-attention in the VLM (Full AK), since the role of the LLM appears negligent so far. Yet, we find is that OPI scores for this setting greatly exceed floor performance, and actually land much closer to the ceiling performance of the No AK baseline. It appears that the language decoder, while passively reading from strong visual representations, can actively “write back” or reconstruct information when presented with weaker visual representations.
@@ 4.4 How Do LLMs Do It?
Prior work shows that different layers in vision encoders specialize in different functions, with early layers extracting low-level features and later layers encoding higher-level semantic information Ghiasi et al. (2022); Dorszewski et al. (2025). Motivated by this, we progressively knock out self-attention in the upper layers of the vision encoder–blocking the last 6, 12, 18, or all 24 layers–and measure how well the language decoder can compensate for the lack in visual contextualization. The results are shown in Figure 5.
Deep Visual Contextualization is Fully Recoverable (by a strong LLM.)
When AK is applied to only the last six layers in LLaVA 13B (last 6 in Figure 5, right) , the identifiability rate remains nearly indistinguishable from the No AK baseline. This finding suggest that the high-level semantic-feature building typically attributed to the last layers of the vision encoder, can be fully reconstructed by the LLM. This result is intuitive in light of studies which establish the isotropy of the semantic spaces learned by even independently trained vision and language models Huh et al. (2024). Yet, to the best of our knowledge, this is the first time this kind of dynamic division of labor between the modules of a VLM has been observed empirically.
The potential for full recovery, moreover, appears to be a function of model size. The results with BakLLaVA (Figure 5, left) show that a considerable gap remains in OPI scores even at six layers of AK in the CLIP encoder (and similarly for LLaVA 7B in Figure 15 in the Appendix.)
@@
As we extend AK in the vision encoder to 12 and then to 18 layers, we observe that the OPI scores converge towards the numbers seen for full (24-layer) attention knockout. There is almost no difference in what the language decoder can recover from visual representations with six layers of early contextualization with zero layers of early contextualization, meaning that some critical low-level, purely visual feature extraction takes place in those first six layers of the encoder, that is beyond the scope LLM decoder.
Nevertheless, it bears repeating that even with full 24-layer AK in the CLIP encoder, the LLM is able to compensate for much of the identifiability of object parts–this was the highlight finding of § 4.3. The analysis above sheds more light on what it the LLM cannot compensate for, which appears to be rooted in shallow-layer visual feature extraction.
The Surge in Identifiability Adaptively Shifts
This is best seen in the results for LLaVA 13B, where the no AK baseline exhibits two clear surges, as discussed in §4.2. In Figure 5 (right), we see that the pattern of the surges is preserved across different levels of attention knockout in the vision encoder. But it shifts from earlier layers to later layers in the decoder, as AK in the vision encoder progresses from the last 6 layers, through 12, 18 and finally 24 layers. This delay reflects the compensatory behavior of the LLM decoder: as there is more missing context in the visual representation, the LLM has to dedicate more layers to recover it, effectively “writing” key information to the visual representation, before reading it back, i.e. before identifying the object part present in a region.
@@ 4.5 Can CLIP Identify Object Parts?
Earlier we saw that different LLaVA variants yield a high OPI score which is largely dependent of the good contextualization of image patches in the CLIP vision encoder. Here, we test whether the CLIP text encoder itself is also able to read this information. Our implementation of the OPI probe is a modification on the standard approach for zero-shot image classification with CLIP.
@@
In zero-shot image classification, CLIP’s text encoder is used to obtain representations for all vocabulary tokens contextualized within a template (“A photo of {token}”.) and the <|endoftext|> token is used to represent each candidate text. Tokens are then ranked based on their similarity to an input image [CLS] token representation. The standard approach for image classification cannot be used for patch-level analysis off-the-shelf, since the [CLS] token pulls information from all patches in the image. So we implement a simple localization method through self-attention blocking to ensure that the [CLS] token represents a specific image region. Concretely, we allow self-attention to be computed as usual up to a given layer, $l$, in the vision encoder, while for all subsequent layers, self-attention is constrained to the [CLS] token and patches of the target region. Unlike the AK interventions used earlier, the procedure described here is not a form of attention ablation, but a method used to focus the representational power of the [CLS] token onto the region of interest within the image. We empirically test different values for $l$ and find that $l=22$ works best (see Figure 9 in the Appendix).
Figure 6 presents CLIP’s OPI scores, including earlier LLaVA 13B results for reference.
@@ 6 Limitations
While our study provides valuable insights into the contextualization mechanisms of vision-language models, it also has some limitations. We use the Object-Part Identification (OPI) task which, despite no direct applied value, serves as an effective diagnostic probe for isolating and analyzing specific model behaviors in a controlled setting. In this analysis, we rely on a single dataset–Pascal Panoptic Parts which we carefully filter further. We selected this dataset primarily due to detailed, structured annotations of objects and their corresponding parts, which are essential for our task at hand. Most of the other segmentation datasets lack this level of granularity, making them unsuitable for studying contextualization in the way we do in this work.

### v2

@@ Abstract
Many vision-language models (VLMs) that prove very effective at a range of multimodal tasks build on CLIP-based vision encoders, which are known to have various limitations. We investigate the hypothesis that the strong language backbone in VLMs compensates for possibly weak visual features by contextualizing or enriching them. Using three CLIP-based VLMs, we perform controlled self-attention ablations on a carefully designed probing task. Our findings show that despite known limitations, CLIP visual representations offer ready-to-read semantic information to the language decoder. However, in scenarios of reduced contextualization in the visual representations, the language decoder can largely compensate for the deficiency and recover performance. This suggests a dynamic division of labor in VLMs and motivates future architectures that offload more visual processing to the language decoder.
@@ 1 Introduction
Vision-language models (VLMs) have made remarkable progress in recent years, with systems like MAGMA (Eichenberg et al., 2021), BLIP-2 (Li et al., 2023a), LLaVA Liu et al. (2023, 2024a), and Prismatic (Karamcheti et al., 2024) demonstrating strong performance for their time on key multimodal benchmarks. A common design choice across these models is the use of a frozen pretrained vision encoder, often based on CLIP Radford et al. (2021), paired with a pre-trained language model, which is finetuned to map visual features into text. [^1] The vision encoder and language decoder are commonly connected with a linear projection or a multi-layer perceptron, which has largely been shown to serve the technical role of mapping between dimensionalities rather than semantic spaces (Schwettmann et al., 2023a; Verma et al., 2024).
@@ 2 Related Work
Zooming in on the vision encoder itself, Gandelsman et al. (2024, 2025) decomposed CLIP’s visual embeddings into sparse text concepts and found that different attention heads specialize in distinct semantic properties (e.g., color, shape, location, etc). Despite CLIP’s strong zero-shot classification performance, various studies have highlighted its limitations, ranging from poor compositional reasoning and limited sensitivity to fine visual distinctions (Wang et al., 2025; Shao et al., 2023; Lewis et al., 2024; Zhang et al., 2024b), to biases related to object quantity and size Abbasi et al. (2025).
@@
Figure 1: Overview of Object Part Identification. Given an object part (e.g., the tail of a cat), localized to a region of image patches through a segmentation mask, a representation is obtained from the VLM for all relevant patches, a probability distribution is induced through LogitLens, and an identifiability rank is extracted from each distribution for the relevant label (tail). The highest rank across patches indicates the overall part identifiability.
Together, these studies tell a nuanced story of the mechanisms behind building visual representations, with their perceived and real limitations, and translating those into text. Our work investigates how visual representations evolve through the layers of the language decoder and to what extent their deficiencies are resolved in the process.
@@ 3 Object Part Identification (OPI)
OPI assumes the availability of patch-level masks which localize object parts in an input image. A single patch can contain multiple parts and a single part can span multiple patches. Probing the VLM consists of: (1) inducing a probability distribution from a VLM for every image patch, (2) extracting the rank of all part labels relevant to a patch, and (3) aggregating the ranks across all patches in the part-relevant. This results in one value per part which indicates how well the model identifies this part in this image.
@@
We derive the necessary annotations from the Pascal Panoptic Parts (Pascal-PP) dataset (de Geus et al., 2021; Meletis et al., 2020) which provides high-quality pixel-level segmentation annotations for 194 part classes spanning 16 object classes, e.g., dog-head, horse-tail, person-leg, etc. We focus on 7 out of the 16 object classes available in Pascal-PP, specifically animal objects. This and other measures taken to reduce noise are outlined in Appendix˜A. After filtering, we sample up to 100 images per class, where available.
Finally, we convert pixel-level segmentation masks into patch-level masks, effectively reducing the resolution of the masks. The final dataset contains 567 unique images and 2840 unique part regions. (see Table˜1 for more details.)
@@
Prior to observing the rank of a label, the probability distribution is processed to aggregate all label aliases, e.g., “leg”, “legs”, “_leg”, “_legs”, under a single token, leg, summing up all their probabilities. [^3] The label aliases are manually defined with reference to the VLM vocabulary. The rank of label $l$ is obtained from the probability distribution $p_{i}$ as follows:
@@ 4.1 Model Architecture and Details
Table 1: Summary of selected object classes with available parts. NN: total number of images per class and $P$: total number of unique part regions per class.
We carry out experiments with LLaVA-1.5 7B and 13B [^4] https://huggingface.co/collections/llava-hf/llava-15-65f762d5b6941db5c2ba07e0 , BakLLaVA 7B [^5] https://huggingface.co/llava-hf/bakLlava-v1-hf and TinyLLaVA [^6] https://huggingface.co/bczhou/tiny-llava-v1-hf (Zhou et al., 2024), all three of which use the CLIP ViT-L/14 image encoder with 24 layers (Radford et al., 2021). As decoder, LLaVA-1.5 7B/13B uses Vicuna 7B/13B (Chiang et al., 2023), BakLLaVA uses Mistral 7B v0.1 (Jiang et al., 2023), and TinyLLaVA uses TinyLLaMA (1.1B) (Zhang et al., 2024a). All language decoders are prompted using their standard template (“USER: <image>”).
This selection of models with different language backbones allows us to robustly study the role of visual input contextualization in LLMs in general. Indeed, our results show consistent trends across all VLMs, with reduced magnitude for the considerably smaller TinyLLaVA model, whose results we show Table˜3 in the Appendix.
@@ 4.2 Establishing a Baseline
Using the method introduced in §3, we obtain OPI scores from all models. The scores, averaged across all parts per object (bird, cat, etc.) and across all objects (all), can be seen in Figure˜2 (left). Average object identifiability scores are also included for reference (hatched bars), as induced from part regions. We observe that the OPI rate is significantly above chance for all three models, at or above 80% on average, and matches the object identifiability. This shows that the VLMs effectively map between visual concepts and their corresponding vocabulary tokens even on the low level of object parts. While it is expected that object information is often mentioned in text, object parts are mentioned far more rarely.
In Figure˜2 (right), we see how identifiability progresses across the layers of the models, as a function of percentage of layers computed. OPI in LLaVA 13B progresses in two steps, with one surge at about 20% and another at 40% after which it nearly plateaus. LLaVA 7B shows a steep ascent up to about 60% of its layers, and BakLLaVA’s progression is stable throughout. Interestingly, all models show a drop in identifiability in the last 2-3 layers (with a subsequent recovery in the last layer for LLaVA 7B only). This LLM behavior might be linked to confidence calibration mechanisms, which aim to prioritize a specific set of output tokens, subject to strong language priors in addition to visual context.
@@ 4.3 Ablating Contextualization
Having established that LLaVA is highly successful in identifying object parts, next we inspect the source of this performance by ablating the contextualization of image patches in the visual encoder and the language decoder through Attention Knockout (AK) (Geva et al., 2023). We block the flow of information between the part regions and the rest of the image, at different layers of LLaVA. This allows us to assess how contextualization plays into OPI across different layers of the encoder and decoder.
@@
Figure˜12 presents the layer-wise identifiability results for all experimental configurations, with a per-object breakdown for LLaVA 13B and averaged results for all three models. Below, each configuration is discussed in turn.
@@
To assess the role of contextualization in the OPI task, we establish a floor of performance by applying attention knockout to both the vision encoder and the language decoder (Full AK). This leads to a sizable drop in OPI scores compared to the baseline (No AK), which limited recovery across the layers of the decoder. Despite this drop, the identifiability rate remains nonzero, indicating that some object parts can be identified to a high degree even in isolation from the object they belong to. In Figure˜4, we show one such case: the leg of a person, whose identifiability drops a bit as a result of the full attention knockout, but stands apart from other parts in how little its identifiability changes. Overall, however, OPI proves to be a highly context-dependent task, as intended.
Trends are similar for BakLLaVA 7B and LLaVA 7B (last column in Figure˜4), with the gap between the upper (No AK) bound and lower (Full AK) bound being smaller in comparison to LLaVA 13B, but still considerable.
@@
Based on the previous two AK configurations, we would expect that blocking self-attention in the vision encoder (AK in CLIP Encoder) should have an almost as detrimental effect as blocking all self-attention in the VLM (Full AK), since the role of the LLM appears negligent so far. Yet, we find that OPI scores for this setting greatly exceed floor performance, and actually land much closer to the ceiling performance of the No AK baseline. It appears that the language decoder, while passively reading from strong visual representations, can actively “write back” or reconstruct information when presented with weaker visual representations.
@@ 4.4 How Do LLMs Do It?
Prior work shows that different layers in vision encoders specialize in different functions, with early layers extracting low-level features and later layers encoding higher-level semantic information Ghiasi et al. (2022); Dorszewski et al. (2025). Motivated by this, we progressively knock out self-attention in the upper layers of the vision encoder–blocking the last 6, 12, 18, or all 24 layers–and measure how well the language decoder can compensate for the lack in visual contextualization. The results are shown in Figure˜5.
Deep Visual Contextualization is Fully Recoverable (by a strong LLM).
When AK is applied to only the last six layers in LLaVA 13B (last 6 in Figure˜5, right), the identifiability rate remains nearly indistinguishable from the No AK baseline. This finding suggests that the high-level semantic-feature building typically attributed to the last layers of the vision encoder can be fully reconstructed by the LLM. This result is intuitive in light of studies which establish the isotropy of the semantic spaces learned by even independently trained vision and language models Huh et al. (2024). Yet, to the best of our knowledge, this is the first time this kind of dynamic division of labor between the modules of a VLM has been observed empirically.
The potential for full recovery, moreover, appears to be a function of model size. The results with BakLLaVA (Figure˜5, left) show that a considerable gap remains in OPI scores even at six layers of AK in the CLIP encoder (and similarly for LLaVA 7B in Figure˜15 in the Appendix.)
@@
As we extend AK in the vision encoder to 12 and then to 18 layers, we observe that the OPI scores converge towards the numbers seen for full (24-layer) attention knockout. There is almost no difference in what the language decoder can recover from visual representations with six layers of early contextualization with zero layers of early contextualization, meaning that some critical low-level, purely visual feature extraction takes place in those first six layers of the encoder, which is beyond the scope of the LLM decoder.
Nevertheless, it bears repeating that even with a full 24-layer AK in the CLIP encoder, the LLM is able to compensate for much of the identifiability of object parts – this was the highlight finding of § 4.3. The analysis above sheds more light on what the LLM cannot compensate for, which appears to be rooted in shallow-layer visual feature extraction.
The Surge in Identifiability Adaptively Shifts.
This is best seen in the results for LLaVA 13B, where the no AK baseline exhibits two clear surges, as discussed in §4.2. In Figure˜5 (right), we see that the pattern of the surges is preserved across different levels of attention knockout in the vision encoder. But it shifts from earlier layers to later layers in the decoder, as AK in the vision encoder progresses from the last 6 layers, through 12, 18 and finally 24 layers. This delay reflects the compensatory behavior of the LLM decoder: as there is more missing context in the visual representation, the LLM has to dedicate more layers to recover it, effectively “writing” key information to the visual representation, before reading it back, i.e., before identifying the object part present in a region.
@@ 4.5 Can CLIP Identify Object Parts?
Earlier we saw that different LLaVA variants yield a high OPI score which is largely dependent on the good contextualization of image patches in the CLIP vision encoder. Here, we test whether the CLIP text encoder itself is also able to read this information. Our implementation of the OPI probe is a modification of the standard approach for zero-shot image classification with CLIP.
@@
In zero-shot image classification, CLIP’s text encoder is used to obtain representations for all vocabulary tokens contextualized within a template (“A photo of {token}”.) and the <|endoftext|> token is used to represent each candidate text. Tokens are then ranked based on their similarity to an input image [CLS] token representation. The standard approach for image classification cannot be used for patch-level analysis off-the-shelf, since the [CLS] token pulls information from all patches in the image. So we implement a simple localization method through self-attention blocking to ensure that the [CLS] token represents a specific image region. Concretely, we allow self-attention to be computed as usual up to a given layer, $l$, in the vision encoder, while for all subsequent layers, self-attention is constrained to the [CLS] token and patches of the target region. Unlike the AK interventions used earlier, the procedure described here is not a form of attention ablation, but a method used to focus the representational power of the [CLS] token onto the region of interest within the image. We empirically test different values for $l$ and find that $l=22$ works best (see Figure˜9 in the Appendix).
Figure˜6 presents CLIP’s OPI scores, including earlier LLaVA 13B results for reference.
@@ 6 Limitations
While our study provides valuable insights into the contextualization mechanisms of vision-language models, it also has some limitations. We use the Object-Part Identification (OPI) task which, despite no direct applied value, serves as an effective diagnostic probe for isolating and analyzing specific model behaviors in a controlled setting. In this analysis, we rely on a single dataset–Pascal Panoptic Parts which we carefully filter further. We selected this dataset primarily due to detailed, structured annotations of objects and their corresponding parts, which are essential for our task at hand. Most of the other segmentation datasets lack this level of granularity, making them unsuitable for studying contextualization in the way we do in this work. Furthermore, our study is limited to LLaVA-style models (Liu et al., 2023, 2024a; Zhou et al., 2024), which pass all patch representations directly to the LLM decoder; our method does not extend to architectures like BLIP-2 Li et al. (2023b), Qwen2-VL (Wang et al., 2024) or Phi-3.5-Vision [^7] https://huggingface.co/microsoft/Phi-3.5-vision-instruct, which instead rely on sampling or pooling of the original patch representations. We leave the exploration of how our probing framework can be adapted to these newer architectures for future work.